# From chimeras to extensive chaos in networks of heterogeneous Kuramoto oscillator populations

### v1

@@ Abstract
Networks of coupled oscillators can exhibit a wide range of complex dynamical behavior, from complete synchronization to chimera and chaotic states. Here we analyze the dynamics of networks of populations of heterogeneous mean-field coupled Kuramoto-Sakaguchi oscillators and show that the instability that leads to chimera states in a simple two-population model also leads to extensive chaos in large networks of coupled populations. Formally, the system consists of a complex network of interacting oscillator populations whose mesoscopic behavior evolves according to the Ott-Antonsen equations. By considering identical parameters across populations, the system contains a manifold of homogeneous solutions where all populations behave identically. Stability analysis of these homogeneous states provided by the Master Stability Function shows that non-trivial dynamics might emerge on a wide region of the parameter space for arbitrary network topologies. As examples, we first revisit the two-population case, and provide a complete bifurcation diagram. Then, we investigate the emergent dynamics in large ring and Erdös-Rényi networks. In both cases, transverse instabilities lead to extensive space-time chaos, i.e., irregular regimes whose complexity scales linearly with the system size. In summary, our work provides a unified analytical framework to understand the emergent dynamics of networks of Kuramoto populations, from chimera states to robust high-dimensional chaos.
@@ I Introduction
The Kuramoto-Sakaguchi modelSakaguchi and Kuramoto (1986) provides a simple yet powerful framework to understanding the collective behavior of complex oscillatory systems, including chimera states. In this model, each oscillator corresponds to a single phase variable which interacts with other units through all-to-all sinusoidal coupling. Remarkably, this mean-field setup allows an exact low-dimensional reduction through the Ott-Antonsen ansatzOtt and Antonsen (2008); Bick et al. (2020). This method provides the collective dynamics of the system as a single differential equation for the complex Kuramoto order parameter  [^1] See also the Watanabe-Strogatz theoryWang and Buzsáki (1996), which provides a complete description for the homogeneous case. For a comprehensive general theory that includes and relates Watanabe-Strogatz and Ott-Antonsen frameworks see Ref. [Cestnik and Pikovsky, 2022]. .
@@
Researchers have employed the Ott-Antonsen equations to unveil the nature of chimera states in systems composed of two identical populations of Kuramoto-Sakaguchi oscillatorsAbrams and Strogatz (2004); Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021). For oscillators with equal natural frequencies, the chimera state always coexist with the homogeneous solution (i.e., full synchronization of the two populations)Abrams et al. (2008). However, for oscillators with distributed frequencies, the chimera state might emerge through a pitchfork bifurcation of the homogeneous solution, thus attracting solutions are always inhomogeneous in wide regions of the parameter spaceLaing (2009a); Kotwal, Jiang, and Abrams (2017). Interestingly, chimera states are not always stable in these regions, leading to the emergence of other regimes including symmetric oscillations, antiphase states, or even low-dimensional chaosLaing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021)  [^2] For studies of non-identical populations (i.e., ensembles governed by different parameter values) see [Montbrió, Kurths, and Blasius, 2004; Martens, Bick, and Panaggio, 2016; Choe, Ri, and Kim, 2016]. .
@@
A natural question to ask is how do these results transduce to large networks of more than two populations. A number of recent studies have addressed this topicMartens (2010a, b); Lee and Krischer (2023a, b); Laing (2023) (see also [Bick et al., 2020] for a thorough review). In particular, Martens [Martens, 2010a, b] studied networks of three populations without disorder, and unveiled the emergence of different types of chimera states. Lee and Krischer [Lee and Krischer, 2023a] extended these studies and showed the emergence of low-dimensional chaos in the three-population model. The same authors also investigated six populations on a ring topology, and showed that chimeras are generally unstable, but connected through attracting heteroclinic cyclesLee and Krischer (2023b). On the other hand, Laing [Laing, 2023] investigated the emergence of chimera states in rings composed of different number of populations with non-local coupling. Interestingly, this work unveils chaotic dynamics appearing through the instability of chimera states, and shows that this irregular regime vanishes as the number of populations in the network increases.
@@
In spite of all this progress, the behavior of identical populations Kuramoto-Sakaguchi populations interacting through complex networks remains an open problem due to the variety of possible topological configurations. In this paper, we take on this challenge, and provide a general unifying framework to study the emergence of spatiotemporal dynamics in these models. We consider networks where the overall external input to each population is normalized across nodes, ensuring thus the existence of homogeneous states. The stability of these trivial regimes to small perturbations can be studied by means of the Master Stability FunctionPecora and Carroll (1998), which provides the growth rate of arbitrary perturbations as a function of the structural connectivity eigenmodes. This method allows to obtain an analytical expression for the dispersion relation of the trivial states.
@@
We show that, for populations with intrinsic disorder, transverse instabilities of the homogeneous state arise in a wide region of the parameter space for arbitrary network topologies. In the two-population model, this transverse instability corresponds to the pitchfork bifurcation leading to chimera statesLaing (2009a); Kotwal, Jiang, and Abrams (2017), whereas in large networks of ring and Erdős-Rényi topologies, the dynamics becomes immediately chaotic with several positive Lyapunov exponents. Numerical simulations reveal this is a case of extensive chaos, a dynamical regime in which the dimension of the chaotic attractor increases linearly with the system size. Extensive chaos, first conjectured by RuelleRuelle (1982), has been widely studied in spatially extended systems (usually under the more general term of "space-time chaos")Xi et al. (2000); Paul et al. (2007), networks of spiking neuronsMonteforte and Wolf (2010); Palmigiano, Engelken, and Wolf (2022), and mean-field modelsTakeuchi et al. (2011) among other systems. Nonetheless, to the best of our knowledge, this phenomena had never been reported in networks of Kuramoto oscillators before.
@@
Overall, our study presents a general analytical approach to study the dynamics of networks of identical Kuramoto-Sakaguchi populations, and provides a direct link between the emergence of chimera states in two populations and the appearance of extensive chaos in large networks.
@@ II A network of Kuramoto-Sakaguchi oscillator populations
Let us consider $N$ populations composed of $M$ Kuramoto-Sakaguchi oscillators each Sakaguchi and Kuramoto (1986); Bick et al. (2020). Each oscillator in a population is described by their phase variable $\phi_{j,\sigma}$ where $j=1,\dots,M$ and $\sigma\in 1,\dots,N$. The phase dynamics are given by
@@
  • We assume $C$ row-normalized, i.e., $\sum_{\kappa=1}^{N}c_{\sigma\kappa}=1\;,$ thus, the total influence received by each oscillator is the same across populations.
  • We assume $C$ is diagonalizable. This condition is always fulfilled if the network is undirected.
Generally we focus in cases where the internal coupling within each population is larger than the coupling between other populations, i.e., $c_{\sigma\sigma}>c_{\sigma\kappa}$ for all $\sigma\neq\kappa$. Nonetheless, this is not a strict requirement.
The natural frequencies of the oscillators $\omega_{j,\sigma}$ are distributed according to a Cauchy (Lorenz) distribution centered at $\overline{\omega}$ and with half-width at half-maximum $\Delta$. Due to the rotational symmetry of the model, we can set $\overline{\omega}=\tan(\alpha)(K\cos(\alpha)-\Delta)$ without loss of generality (see Appendix C), a choice that will be justified in the next section.
@@
This equation provides the mean-field dynamics of each population in the network, and its attractiveness for $\Delta>0$ has been recently provenCestnik and Pikovsky (2022). Therefore, Eq. (3) is going to be our main object of study in the rest of the article.
@@ III Analysis of homogeneous states and their stability
In order to explore the dynamics of (3), first we study its homogeneous manifold, i.e., all the states in which the $N$ populations evolve with exactly the same dynamics. By imposing $Z_{\sigma}=Z$ $\forall\sigma$, into Eq. (3) and exploiting the fact that $C$ is row-normalized, we obtain
@@
which corresponds to the mean-field equation for a single, isolated population of Kuramoto-Sakaguchi oscillators Ott and Antonsen (2008); Bick et al. (2020). The dynamics of this system are well studied: if $K<K_{c}:=\frac{2\Delta}{\cos(\alpha)}$ the system converges to the incoherent state $Z^{(0)}:=0$. At $K=K_{c}$ the system undergoes a supercritical Hopf bifurcation (H0 in the figures) giving rise to a synchronized or coherent state $Z:=Re^{i\Omega t}$ with
@@
These equations show that our choice of the mean frequency $\overline{\omega}=\tan(\alpha)(K\cos(\alpha)-\Delta)$ defines a co-rotating frame in which the homogeneous coherent solution is a steady state rather than a limit-cycle (i.e., $\Omega=0$). This simplifies the following analytical calculations.
Next, we focus on the stability of the coherent homogeneous state $Z$ given by Eq. (5). The same analysis for the incoherent homogeneous solution $Z^{(0)}$ is provided in Appendix E, where we show that no instabilities occur below the Kuramoto synchronization transition $H_{0}$. Let us consider a small perturbation of the coherent homogeneous state, $Z_{\sigma}=Z+z_{\sigma}$. Linearizing Eq. (3) we obtain
@@
In order to lighten notation, in the following steps we express complex quantities in a vectorial format: $z=(x,y)^{T}$ where $x$ and $y$ are the real and imaginary parts of $z$ respectively. Then, Eq. (6) can be expressed as
$\dot{z}_{\sigma}=Az_{\sigma}+B\sum_{\kappa=1}^{M}c_{\sigma\kappa}z_{\kappa}$ (7)
@@
are obtained by inserting the values of $\overline{\omega}$ and $R$ derived from Eq. (5) into Eq. (6) and separating real and imaginary parts.
Next, we apply the Master Stability Formalism Pecora and Carroll (1998) in order to simplify the analysis of Eq. (7)333 See Refs. [Arenas et al., 2008; Porter and Gleeson, 2016; Ashwin, Coombes, and Nicks, 2016] for introductory reviews on the MSF formalism. Although most studies on the MSF rely on diffusive coupling or variations of it via a Laplacian matrix (see, e.g., [Nakao, 2014]), the method is applicable to any setup having a homogeneous solution (see, e.g., [Clusella et al., 2023]). . This method consists on decomposing the perturbation $z_{\sigma}$ on the basis provided by the diagonalization of the connectivity matrix $C$. We denote the eigenvalues of the connectivity matrix as $\Lambda_{1}\geq\Lambda_{2}\geq\cdots\geq\Lambda_{N}$, and the corresponding eigenvectors as $\bm{\Psi}^{(k)}=(\Psi_{1}^{(k)},\dots,\Psi_{N}^{(k)})^{T}$ with $k=1,\dots,N$. Therefore
@@
Then, we express the perturbation $z_{\sigma}$ on the basis composed of $\bm{\Psi}_{k}$ for $k=1,\dots,N$ as
$z_{\sigma}=\sum_{k=1}^{N}u_{k}\otimes\Psi_{\sigma}^{(k)}$ (11)
@@
$=\sum_{k=1}^{N}\left[Au_{k}\otimes\Psi_{\sigma}^{(k)}+Bu_{k}\otimes\sum_{\kappa=1}^{N}c_{\sigma\kappa}\Psi_{\kappa}^{(k)}\right]$
$=\sum_{k=1}^{N}\left[Au_{k}\otimes\Psi_{\sigma}^{(k)}+\Lambda_{k}Bu_{k}\otimes\Psi_{\sigma}^{(k)}\right]$
$=\sum_{k=1}^{N}\left[A+\Lambda_{k}B\right]u_{k}\otimes\Psi_{\sigma}^{(k)}.$
Then, since the eigenvectors $\bm{\Psi}^{(k)}$ compose a basis of $\mathbb{R}^{N}$, linear independence provides
$\dot{u}_{k}=\left[A+\Lambda_{k}B\right]u_{k}\;.$ (13)
@@
The eigenvalues $\lambda_{k}^{(\pm)}$ depend explicitly on the parameters of the system ($\Delta$, $\alpha$, and $K$), as well as the structural eigenvalues $\Lambda_{k}$. Using the row-normalization of $C$ and the Gershgorin circle theoremWilkinson (1967) we can assert that $|\Lambda_{k}|\geq 1$. Moreover, the largest structural eigenvalue is $\Lambda_{1}=1$, and the corresponding eigenvector corresponds to a uniform perturbation, i.e.,
@@
For the case $\Delta=0$ the expression of the system eigenvalues simplify to $\lambda^{(+)}=K\cos(\alpha)(\Lambda-1)$ and $\lambda^{(-)}=-K\cos(\alpha)$. Since the coherent state emerges for $K\cos(\alpha)>2\Delta=0$, this proves that, in networks without disorder, the homogeneous state is always stable, consistent with the fact that chimera states in these systems always coexist with the full synchronized solution Abrams et al. (2008).
@@
For $\Delta>0$, however, the scenario changes. In this case it is possible to perform parameter a reduction (see Appendix C), thus we set $\Delta=1$ without loss of generality, and analyze the system for different values of $K$ and $\alpha$. Fig. 1(a) shows $\operatorname{Re}[\lambda^{(\pm)}]$ as a function of $\Lambda$ for different parameter sets. For the selected values, the dispersion relation displays a positive region, indicating the existence of transverse instabilities, which always emerge through real eigenvalues $\lambda^{(+)}\in\mathbb{R}$ 444 The homogeneous synchronized solution emerges only for $\Delta-K\cos(\alpha)/2<0$, which leads to $\Delta+K\cos(\alpha)\left(\frac{\Lambda}{2}-1\right)<0$ in Eq. (14) (since $\Delta\geq 0$ and $-1\leq\Lambda\leq 1$). Therefore, eigenvalues with positive real part in Eq. (14) always correspond to real eigenvalues. Complex conjugate eigenvalues might exist, but they correspond to stable directions. . Overall, the shape of the dispersion relation resembles that of the Benjamin-Feir instability in the complex Ginzburg-Landau equation Kuramoto (1984).
@@
Figure 1: Transverse instabilities of homogeneous states in networks of Kuramoto-Sakaguchi populations Eq. (3). (a) Real part of the eigenvalues characterizing the stability of the homogeneous solution $\lambda^{(\pm)}$ as a function of the structural matrix eigenvalues $\Lambda$. Thick and thin curves correspond to $\lambda^{(+)}$ and $\lambda^{(-)}$ as given by the Master Stability Function in Eq. (14) respectively. Green, orange, and purple curves correspond to $(K,\alpha)=(5,1)$, $(10,1.2)$, and $(15,1.4)$. (b) Color-shaded regions show the regions of transverse instability in the $(K,\Lambda)$ plane, with solid lines showing the loci of the bifurcation ($\text{Re}[\lambda^{(+)}]=0$ in Eq. (14)). Dotted vertical lines indicate the Kuramoto synchronization transition $K_{c}=2\Delta/\cos(\alpha)$ at which the coherent homogeneous solution emerges. Results shown for $\alpha=1$ (red), 1.2 (black), and 1.4 (blue).
@@
Setting $\operatorname{Re}[\lambda^{(+)}]=0$ in Eq (14) determines the regions where transverse instabilities of the homogeneous state arise in the parameter space. Figure 1(b) shows such regions for three different values of $\alpha$ in the $(K,\Lambda)$-plane. A small region of instabilities arise already for $\alpha=1$, and widens as $\alpha\to\pi/2$, the critical value at which the homogeneous coherent solution vanishes through $H_{0}$. This is consistent with previous findings showing that chimera and other symmetry-breaking states emerge only for $\alpha$ close but below $\pi/2$Abrams and Strogatz (2004); Abrams et al. (2008); Kotwal, Jiang, and Abrams (2017). In the next sections we review the dynamics that emerge in this region in the two-population model, ring networks, and Erdős-Rényi connectivites. For this analysis we fix $\alpha=1.2$, a value not that close to the $H_{0}$ bifurcation, and still displaying a large instability region.

### v2

@@ Abstract
Populations of coupled oscillators can exhibit a wide range of complex dynamical behavior, from complete synchronization to chimera and chaotic states. We can thus expect complex dynamics to arise in networks of such populations. Here we analyze the dynamics of networks of populations of heterogeneous mean-field coupled Kuramoto-Sakaguchi oscillators, and show that the instability that leads to chimera states in a simple two-population model also leads to extensive chaos in large networks of coupled populations. Formally, the system consists of a complex network of oscillator populations whose mesoscopic behavior evolves according to the Ott-Antonsen equations. By considering identical parameters across populations, the system contains a manifold of homogeneous solutions where all populations behave identically. Stability analysis of these homogeneous states provided by the master stability function formalism shows that non-trivial dynamics might emerge on a wide region of the parameter space for arbitrary network topologies. As examples, we first revisit the two-population case, and provide a complete bifurcation diagram. Then, we investigate the emergent dynamics in large ring and Erdös-Rényi networks. In both cases, transverse instabilities lead to extensive space-time chaos, i.e., irregular regimes whose complexity scales linearly with the system size. Our work provides a unified analytical framework to understand the emergent dynamics of networks of oscillator populations, from chimera states to robust high-dimensional chaos.
@@
Many physical systems are composed of large numbers of interacting oscillatory units. Such complex systems typically display synchronization patterns that vary in complexity, depending on the oscillators’ intrinsic dynamics and the nature of their interactions. For instance, the Kuramoto-Sakaguchi model provides a simple framework in which a population of globally coupled phase oscillators transitions from incoherence to synchronization as the strength of the coupling is varied. Interestingly, when coupling two populations of Kuramoto-Sakaguchi oscillators other collective dynamics may emerge, in which one population displays a higher degree of synchronization than the other. These regimes are an example of chimera states, a phenomena that has been studied in a number of theoretical setups and experimental works. Here, we extend the analysis of these symmetry-breaking transitions to the case of $N$ oscillator populations interacting through arbitrary complex networks. In particular, for ring and random topologies, we reveal the existence of highly chaotic states whose complexity increases linearly with the system size, a situation known as extensive chaos.
@@ I Introduction
The Kuramoto-Sakaguchi modelSakaguchi and Kuramoto (1986) provides a simple yet powerful framework to understand the collective behavior of complex oscillatory systems, including chimera states. In this model, each oscillator corresponds to a single phase variable that interacts with the other units through an all-to-all sinusoidal coupling. Remarkably, this mean-field setup allows an exact low-dimensional reduction through the Ott-Antonsen ansatzOtt and Antonsen (2008); Bick et al. (2020). This method provides the collective dynamics of the system as a single differential equation for the complex Kuramoto order parameter  [^1] See also the Watanabe-Strogatz theoryWang and Buzsáki (1996), which provides a complete description for the homogeneous case. For a comprehensive general theory that includes and relates Watanabe-Strogatz and Ott-Antonsen frameworks see Ref. [Cestnik and Pikovsky, 2022]. .
@@
Researchers have employed the Ott-Antonsen equations to unveil the nature of chimera states in systems composed of two identical populations of Kuramoto-Sakaguchi oscillatorsAbrams et al. (2008); Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021). For oscillators with equal natural frequencies, the chimera state always coexists with the homogeneous solution (i.e., full synchronization of the two populations)Abrams et al. (2008). However, for oscillators with distributed frequencies, the chimera state might emerge through a pitchfork bifurcation of the homogeneous solution, thus attracting solutions are always inhomogeneous in wide regions of the parameter spaceLaing (2009a); Kotwal, Jiang, and Abrams (2017). Interestingly, chimera states are not always stable in these regions, leading to the emergence of other regimes including symmetric oscillations, antiphase states, or even low-dimensional chaosLaing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021)  [^2] For studies of non-identical populations (i.e., ensembles governed by different parameter values) see Refs.[Montbrió, Kurths, and Blasius, 2004; Martens, Bick, and Panaggio, 2016; Choe, Ri, and Kim, 2016]..
@@
A natural question to ask is how do these results transduce to large networks of more than two populations. A number of recent studies have addressed this topicMartens (2010a, b); Lee and Krischer (2023a, b); Laing (2023) (see also [Bick et al., 2020] for a thorough review). In particular, MartensMartens (2010a, b) studied networks of three populations without disorder, and unveiled the emergence of different types of chimera states. Lee and Krischer Lee and Krischer (2023a) extended these studies and showed the emergence of low-dimensional chaos in the three-population model. The same authors also investigated six populations on a ring topology, and showed that chimeras are generally unstable, but connected through attracting heteroclinic cyclesLee and Krischer (2023b). On the other hand, Laing Laing (2023) investigated the emergence of chimera states in rings composed of different number of populations with non-local coupling. Interestingly, this work unveils chaotic dynamics appearing through the instability of chimera states, and shows that this irregular regime vanishes as the number of populations in the network increases.
@@
In spite of all this progress, the behavior of identical populations of Kuramoto-Sakaguchi populations interacting through complex networks remains an open problem, due to the variety of possible topological configurations. In this paper we take on this challenge, and provide a general unifying framework to study the emergence of spatiotemporal dynamics in these models. We consider networks where the overall external input to each population is normalized across nodes, thus ensuring the existence of homogeneous states. The stability of these trivial regimes to small perturbations can be studied by means of the master stability functionPecora and Carroll (1998), which provides the growth rate of arbitrary perturbations as a function of the eigenmodes of the structural connectivity matrix. This method allows to obtain an analytical expression for the dispersion relation of the homogeneous states.
@@
We show that, for populations with intrinsic disorder, transverse instabilities arise in a wide region of the parameter space for arbitrary network topologies. In the case of two populations, such transverse instability corresponds to the pitchfork bifurcation known to lead to chimera statesLaing (2009a); Kotwal, Jiang, and Abrams (2017), whereas in large networks of ring and Erdős-Rényi topologies, the dynamics becomes immediately chaotic with several positive Lyapunov exponents. Numerical simulations reveal that this is a case of extensive chaos, a dynamical regime in which the dimension of the chaotic attractor increases linearly with the system size. Extensive chaos, first conjectured by RuelleRuelle (1982), has been widely studied in spatially extended systems (usually under the more general term of "space-time chaos")Manneville (1985); Livi, Politi, and Ruffo (1986); Keefe (1989); O’Hern, Egolf, and Greenside (1996); Egolf et al. (2000); Xi et al. (2000); Paul et al. (2007), networks of spiking neuronsMonteforte and Wolf (2010); Palmigiano, Engelken, and Wolf (2022), and mean-field modelsTakeuchi et al. (2011), among other systems. Nonetheless, to the best of our knowledge, this phenomenon had never been reported in networks of Kuramoto oscillators before.
@@
Overall, our study presents a general analytical approach to study the dynamics of networks of identical Kuramoto-Sakaguchi populations, and provides a direct link between the emergence of chimera states in two-population models and the appearance of extensive chaos in large networks of such populations.
@@ II A network of Kuramoto-Sakaguchi oscillator populations
Let us consider $N$ populations composed of $M$ Kuramoto-Sakaguchi oscillators each Sakaguchi and Kuramoto (1986); Bick et al. (2020). Each oscillator in a population is described by their phase variable $\phi_{j,\sigma}$ where $j\in\{1,\dots,M\}$ and $\sigma\in\{1,\dots,N\}$. The phase dynamics are given by
@@
  • We assume that $C$ is row-normalized, i.e., $\sum_{\kappa=1}^{N}c_{\sigma\kappa}=1\;,$ thus, the total influence received by each oscillator is the same across populations.
  • We assume that $C$ is diagonalizable. This condition is always fulfilled if the network is undirected.
Generally, we focus on cases where the internal coupling within each population is larger than the coupling between other populations, i.e., $c_{\sigma\sigma}>c_{\sigma\kappa}$ for all $\sigma\neq\kappa$. Nonetheless, this is not a strict requirement.
The natural frequencies of the oscillators $\omega_{j,\sigma}$ are distributed according to a Cauchy (Lorenz) distribution centered at $\overline{\omega}$ and with half-width at half-maximum $\Delta$. Due to the rotational symmetry of the model, we can set $\overline{\omega}=\tan(\alpha)(K\cos(\alpha)-\Delta)$ without loss of generality, a choice that will be justified in the next section.
@@
This equation provides the mean-field dynamics of each population in the network assuming the phases follow a circular Cauchy distribution (Poisson kernel). Moreover, it has been recently shown that the this ansatz is a global attractor of the Kuramoto-Sakaguchi model for $\Delta>0$Cestnik and Pikovsky (2022). Therefore, Eq. (3) is going to be our main object of study for the rest of the article.
@@ III Analysis of homogeneous states and their stability
In order to explore the dynamics of Eq. (3), first we study its homogeneous manifold, i.e., all the states in which the $N$ populations evolve with exactly the same dynamics. By imposing $Z_{\sigma}=Z$ $\forall\sigma$ into Eq. (3), and exploiting the fact that $C$ is row-normalized, we obtain
@@
which corresponds to the mean-field equation for a single, isolated population of Kuramoto-Sakaguchi oscillators Ott and Antonsen (2008); Bick et al. (2020). The dynamics of this system are well studied: if $K<K_{c}:=\frac{2\Delta}{\cos(\alpha)}$ the system converges to the incoherent state $Z:=0$. At $K=K_{c}$ the system undergoes a supercritical Hopf bifurcation (H0 in the figures) giving rise to a synchronized or coherent state, which takes the form $Z:=Re^{i\Omega t}$, with
@@
These equations show that our choice of the mean frequency $\overline{\omega}=\tan(\alpha)(K\cos(\alpha)-\Delta)$ defines a co-rotating frame in which the homogeneous coherent solution is a steady state rather than a limit-cycle (i.e., $\Omega=0$). This simplifies the analytical calculations that follow.
Next we focus on the stability of the coherent homogeneous state $Z>0$ given by Eq. (5). The same analysis for the incoherent homogeneous solution $Z=0$ is provided in Appendix E, where we show that no instabilities occur below the Kuramoto synchronization transition $\text{H}_{0}$. Let us consider a small perturbation of the coherent homogeneous state, $Z_{\sigma}=Z+z_{\sigma}$. Linearizing Eq. (3) we obtain
@@
In order to simplify the notation, in the following steps we express complex quantities in vectorial form: $z=(x,y)^{T}$ where $x$ and $y$ are the real and imaginary parts of $z$, respectively. Then, Eq. (6) can be expressed as
$\dot{z}_{\sigma}=Az_{\sigma}+B\sum_{\kappa=1}^{M}c_{\sigma\kappa}z_{\kappa},$ (7)
@@
These expressions for $A$ and $B$ are obtained by inserting the values of $\overline{\omega}$ and $R$ derived from Eq. (5) into Eq. (6), and separating real and imaginary parts.
Next we apply the master stability formalism (MSF) Pecora and Carroll (1998), in order to simplify the analysis of Eq. (7)333 See Refs. [Arenas et al., 2008; Porter and Gleeson, 2016; Ashwin, Coombes, and Nicks, 2016] for introductory reviews on the MSF formalism. Although most studies on the MSF rely on diffusive coupling or variations of it via a Laplacian matrix (see, e.g., [Nakao, 2014]), the method is applicable to any setup having a homogeneous solution (see, e.g., [Clusella et al., 2023]). . This method consists on decomposing the perturbation $z_{\sigma}$ on the basis provided by the diagonalization of the connectivity matrix $C$. We denote the eigenvalues of the connectivity matrix as $\Lambda_{1}\geq\Lambda_{2}\geq\cdots\geq\Lambda_{N}$, and the corresponding eigenvectors as $\bm{\Psi}^{(k)}=(\Psi_{1}^{(k)},\dots,\Psi_{N}^{(k)})^{T}$ with $k=1,\dots,N$. Therefore
@@
We now express the perturbation $z_{\sigma}$ on the basis $\bm{\Psi}^{(k)}$ as
$z_{\sigma}=\sum_{k=1}^{N}u_{k}\otimes\Psi_{\sigma}^{(k)},$ (11)
@@
$=\sum_{k=1}^{N}\left(Au_{k}\otimes\Psi_{\sigma}^{(k)}+Bu_{k}\otimes\sum_{\kappa=1}^{N}c_{\sigma\kappa}\Psi_{\kappa}^{(k)}\right)$
$=\sum_{k=1}^{N}\left(Au_{k}\otimes\Psi_{\sigma}^{(k)}+\Lambda_{k}Bu_{k}\otimes\Psi_{\sigma}^{(k)}\right)$
$=\sum_{k=1}^{N}\left(A+\Lambda_{k}B\right)u_{k}\otimes\Psi_{\sigma}^{(k)}.$
Since the eigenvectors $\bm{\Psi}^{(k)}$ are a basis of $\mathbb{R}^{N}$, linear independence provides
$\dot{u}_{k}=\left(A+\Lambda_{k}B\right)u_{k}\;.$ (13)
@@
The eigenvalues $\lambda_{k}^{(\pm)}$ depend explicitly on the parameters of the system ($\Delta$, $\alpha$, and $K$), as well as on the structural eigenvalues $\Lambda_{k}$. Using the row-normalization of $C$ and the Gershgorin circle theoremWilkinson (1967) we can assert that $|\Lambda_{k}|\leq 1$. Moreover, the largest structural eigenvalue is $\Lambda_{1}=1$, and the corresponding eigenvector corresponds to a uniform perturbation, i.e.,
@@
For the case $\Delta=0$, the expression of the system eigenvalues simplifies to $\lambda_{k}^{(+)}=K\cos(\alpha)(\Lambda_{k}-1)$ and $\lambda^{(-)}_{k}=-K\cos(\alpha)$. Since the coherent state emerges for $K\cos(\alpha)>2\Delta=0$, this proves that, in networks without disorder, the homogeneous state is always stable, consistent with the fact that chimera states in these systems always coexist with the fully synchronized solution Abrams et al. (2008).
@@
For $\Delta>0$, however, the scenario changes. In this case it is possible to perform a parameter reduction (see Appendix C). Setting $\Delta=1$ without loss of generality, we analyze the system for different values of $K$ and $\alpha$. Figure 1(a) shows $\operatorname{Re}[\lambda^{(\pm)}]$ as a function of $\Lambda$ for different parameter sets. For the selected values, the dispersion relation displays a positive region, indicating the existence of transverse instabilities, which always emerge through real eigenvalues $\lambda^{(+)}\in\mathbb{R}$ 444 The homogeneous synchronized solution emerges only for $\Delta-K\cos(\alpha)/2<0$, which leads to $\Delta+K\cos(\alpha)\left(\frac{\Lambda}{2}-1\right)<0$ in Eq. (14) (since $\Delta\geq 0$ and $-1\leq\Lambda\leq 1$). Therefore, eigenvalues with positive real part in Eq. (14) always correspond to real eigenvalues. Complex conjugate eigenvalues might exist, but they correspond to stable directions. . Overall, the shape of the dispersion relation resembles that of the Benjamin-Feir instability in the complex Ginzburg-Landau equation Kuramoto (1984).
@@
Figure 1: Transverse instabilities of homogeneous states in networks of Kuramoto-Sakaguchi populations Eq. (3). (a) Real part of the eigenvalues (Floquet exponents) characterizing the stability of the homogeneous solution $\lambda^{(\pm)}$ as a function of the structural matrix eigenvalues $\Lambda$. Thick and thin curves correspond to $\lambda^{(+)}$ and $\lambda^{(-)}$ as given by the master stability function in Eq. (14) respectively. Green, orange, and purple curves correspond to $(K,\alpha)=(5,1)$, $(10,1.2)$, and $(15,1.4)$. (b) Color-shaded regions show the regions of transverse instabilities in the $(K,\Lambda)$ plane, with solid lines showing the loci of the bifurcation ($\text{Re}[\lambda^{(+)}]=0$ in Eq. (14)). Dotted vertical lines indicate the Kuramoto synchronization transition $\text{H}_{0}$ at which the coherent homogeneous solution emerges. Results shown for $\alpha=1$ (red), 1.2 (black), and 1.4 (blue).
@@
Setting $\operatorname{Re}[\lambda^{(+)}]=0$ in Eq. (14) determines the regions where transverse instabilities of the homogeneous state arise in parameter space. Figure 1(b) shows such regions for three different values of $\alpha$ in the $(K,\Lambda)$-plane. A small region of instabilities arise already for $\alpha=1$, and widens as $\alpha\to\pi/2$, the critical value at which the homogeneous coherent solution vanishes through H0. This is consistent with previous findings showing that chimera and other symmetry-breaking states emerge only for $\alpha$ close to, but below, $\pi/2$Abrams and Strogatz (2004); Abrams et al. (2008); Kotwal, Jiang, and Abrams (2017). In the next sections we review the dynamics that emerge in this region in the two-population model, ring networks, and Erdős-Rényi connectivites. For this analysis we fix $\alpha=1.2$, a value not that close to the H0 bifurcation, and still displaying a large instability region.